# Cosmology with Binary Neutron Stars: Does the Redshift Evolution of the Mass Function Matter?

### v1

@@ 5 Data Release and Software
The data used in this work is available on Zenodo under an open-source Creative Commons Attribution license at https://zenodo.org/records/14031505 (catalog 10.5281/zenodo.14031505). The corresponding codes are available here: https://github.com/SoumendraRoy/RedevolBNS.
@@
The authors thank Jose María Ezquiaga, Max Isi, Claire Lamman, Utkarsh Mali, Suvodip Mukherjee, and Aditya Vijaykumar for valuable discussions and their contributions in shaping this manuscript. SKR thanks the Center for Computational Astrophysics at the Flatiron Institute for hospitality while this research was carried out. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. AR acknowledges support from the National Science Foundation award PHY-2207728. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. The construction and operation of KAGRA are funded by Ministry of Education, Culture, Sports, Science and Technology (MEXT), and Japan Society for the Promotion of Science (JSPS), National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS) and the Ministry of Science and Technology (MoST) in Taiwan. This paper carries LIGO document number LIGO-P2400446.
Appendix A Comparison of Astrophysical Observations with the BNS Population in COMPAS
Although the BNS mass function is roughly redshift-independent in our models, its shape varies significantly depending on binary physics choices and initial conditions (see, for example, Fig. 1 in van Son et al. (in prep) and the primary-secondary mass corner plot for simulation variations in Section 2: link). Here, we compare the mass function for the fiducial COMPAS setup with observed astrophysical neutron star mass measurements. We note that this fiducial population aligns with the galactic binary pulsar population inferred by Farrow et al. (2019) and with binary neutron star merger populations observed in GW events (Abbott et al., 2023b).
@@ Appendix B Gaussian Mixture Model Fitting of the True Binary Neutron Star Population
Figure 5: The corner plots show the COMPAS populations of masses and redshifts for both the fiducial setup and the Mandel-Müller kick prescription for remnant mass, along with the GMM fit of these populations. Code to generate these plots is available: here.

### v2

@@ 5 Data Release and Software
The data used in this work is available on Zenodo: https://zenodo.org/records/14704635 (catalog 10.5281/zenodo.14704635). The corresponding codes are available here: https://github.com/SoumendraRoy/RedevolBNS.
@@
The authors thank Jose María Ezquiaga, Max Isi, Claire Lamman, Utkarsh Mali, Ilya Mandel, Suvodip Mukherjee, and Aditya Vijaykumar for valuable discussions and their contributions in shaping this manuscript. SKR thanks the Center for Computational Astrophysics at the Flatiron Institute for hospitality while this research was carried out. The computations in this work were, in part, run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation. AR acknowledges support from the National Science Foundation award PHY-2207728. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. The construction and operation of KAGRA are funded by Ministry of Education, Culture, Sports, Science and Technology (MEXT), and Japan Society for the Promotion of Science (JSPS), National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS) and the Ministry of Science and Technology (MoST) in Taiwan. This paper carries LIGO document number LIGO-P2400446.
Appendix A Comparison of Astrophysical Observations with the Redshift Marginalized BNS Mass Population in COMPAS
Although the BNS mass function is roughly redshift-independent in our models, its shape varies significantly depending on binary physics choices and initial conditions (see, for example, Fig. 1 in and the primary-secondary mass corner plot for simulation variations in Section 2: link). Here, we compare the mass function for the fiducial COMPAS setup with observed astrophysical neutron star mass measurements. We note that this fiducial population aligns with the galactic binary pulsar population inferred by Farrow et al. (2019) and with binary neutron star merger populations observed in GW events (Abbott et al., 2023b).
@@ Appendix B Gaussian Mixture Model Fitting of the True Binary Neutron Star Population
Figure 4: The corner plots show the COMPAS populations of masses and redshifts for both the fiducial setup and the Mandel-Müller kick prescription for remnant mass, along with the GMM fit of these populations. In the upper left plot of the fiducial population, fewer than $10$ events following a northwest-to-southeast path ($\log q\approx-0.7$ to $-0.5$) are not captured by the GMM. However, since the chirp mass of these events still falls within the GMM-covered population-and chirp mass is the most precisely measured parameter in cosmological parameter estimation-there is no resulting bias in the cosmological parameters. Code to generate these plots is available: here.